OpenStudy (anonymous):

help me solve this? descriptions are loved :3

OpenStudy (anonymous):

\[(\sqrt{7}-\sqrt{2})^{2}\]

OpenStudy (anonymous):

using the fact that:\[(a-b)^2=a^2-2ab+b^2 \]we can calculate your problem as:\[(\sqrt{7}-\sqrt{2})^2=(\sqrt{7})^2-2(\sqrt{7})(\sqrt{2})+(\sqrt{2})^2\]\[=7-2\sqrt{7\cdot2}+2=9-2\sqrt{14}\]

OpenStudy (anonymous):

Ok thanks :)